OpenStudy (anonymous):

The of a rectangle is four times its width. If the perimeter is at most 130 centimeters, what is the greatest possible value for the width?

OpenStudy (anonymous):

The ___ of a rectangle is four time its width. what is missing there?

OpenStudy (anonymous):

Well no answer. If it's supposed to be 'length' then it's not supposed to be very complicated \[ Width = x \\ Length = 4x \\ 4x + 4x + x + x \le 130 \\ 10x \le 130 \\ x \le 13 \] so x is at most 13